OpenStudy (anonymous):

2u sqrt(12uw^2) -w sqrt(75u^3) simplify the radical expression

OpenStudy (campbell_st):

well simplifying the radicals look at the terms inside the radicals are show then and find the square factors. \[2u \times \sqrt{4w^2 \times 3u} - u \sqrt{25u^2 \times 3u}\] which simplifies to \[2u \times 2w \sqrt{3u} - 5u \sqrt{3u}\] you can finish it from here